A mixture of argon and xenon gases, in a 8.85 L flask at 44 °C, contains 17.9 grams of argon and 22.6 grams of xenon. The partial pressure of xenon in the flask is____ atm and the total pressure in the flask is ____ atm.


1
Expert's answer
2021-04-25T03:50:48-0400

The total pressure (P) can be calculated using ideal gas equation: PV = nRT,

 

for the above problem n is the total number of moles, thus

 

n =PV/RT

 

V = 8.85 L, T = 44oC + 273 = 317 K, R = 0.0821 L.atm/mol.K (it is a constant).

Molar Mass of argon = 39.948

= 17.9/39.948 = 0.448M

Molar Mass of Xenon = 131.293

= 22.6/131.293 = 0.172M

0.448 +0.172 = 0.62M

PV = n×RT

8.85P = 0.62 × 0.0821 ×317

8.85p= 16.136

total Pressure = 1.82atm

Partial pressure of Xenon

= 0.172/0.62 × 1.82

= 0.5atm